# Calibrating the Confidence of Large Language Models by Eliciting Fidelity

### v1

@@ Abstract
Large language models optimized with techniques like RLHF have achieved good alignment in being helpful and harmless. However, post-alignment, these language models often exhibit overconfidence, where the expressed confidence does not accurately calibrate with their correctness rate. In this paper, we decompose the language model confidence into the Uncertainty about the question and the Fidelity to the answer generated by language models. Then, we propose a plug-and-play method to estimate the confidence of language models. Our method has shown good calibration performance by conducting experiments with 6 RLHF-LMs on four MCQA datasets. Moreover, we propose two novel metrics, IPR and CE, to evaluate the calibration of the model, and we have conducted a detailed discussion on Truly Well-Calibrated Confidence. Our method could serve as a strong baseline, and we hope that this work will provide some insights into the model confidence calibration.
@@
Mozhi Zhang1thanks: Work done during internship at Meituan., Mianqiu Huang1, Rundong Shi2, Linsen Guo 2, Chong Peng 2, Peng Yan 2, Yaqian Zhou1 Xipeng Qiu1thanks: Corresponding author. 1School of Computer Science, Fudan University 2 Meituan {mzzhang22, mqhuang23}@m.fudan.edu.cn {shirundong}@meituan.com {zhouyaqian, xpqiu}@fudan.edu.cn
@@ 1 Introduction
Large language models (LLMs) acquire vast world knowledge and demonstrate powerful capabilities through pre-training Brown et al. (2020); OpenAI (2023); Bubeck et al. (2023). With technologies like RLHF Ouyang et al. (2022) and RLAIF Bai et al. (2022); Lee et al. (2023), large language models can become more helpful and harmless to align with human preferences Askell et al. (2021). However, how to build a more honest system has not yet been fully discussed. An honest model should have a certain understanding of the boundary of its knowledge, that is, knowing what it does not know Yin et al. (2023); Yang et al. (2023b). A plausible method is utilizing the calibrated confidence to estimate the knowledge boundary of language models. For pre-trained language models, the per-token logit can already be considered a well-calibrated confidence score, which implies that pre-trained language models (mostly) know what they know Kadavath et al. (2022).
@@
Figure 1: The reliability diagram of our method. In four different MCQA datasets, our method has demonstrated good calibration effects. The experimental data is derived from GPT-3.5-Turbo.
@@
As shown in Figure 2, by replacing the model’s answer with “All other options are wrong.”, we can assess whether the model had high fidelity to its previously given answer. Inspired by this phenomenon, we decompose the language model confidence into two dimensions: the Uncertainty about the question and the Fidelity to the answer generated by language models. First, if the answers provided by language model are consistent under multiple samplings, it indicates that language model has lower uncertainty regarding that question. Thus, we could utilize the information entropy of the frequency distribution of sampled answers to calculate the model’s uncertainty about a question. Second, we design a novel method to estimate the model’s fidelity to each of its sampled answers. Last, the uncertainty regarding question $\mathcal{Q}$ and the fidelity to the answer $a_{i}$ together determine the model’s confidence. As shown in Figure 1, our proposed method achieved good calibration across different MCQA datasets. Meanwhile, our approach does not require knowledge of the model’s per-token log-probability, making it broadly applicable to various commercialized RLHF-LMs, which do not provide the per-token log-probability.
To have a closer look at the calibration of model confidence, we propose two novel metrics for evaluating calibration: 1) Inverse Pair Ratio (IPR), which is the proportion of inverse pairs in the Reliability Diagram. This metric could reflect whether the model is well-calibrated from the perspective of the monotonicity of the Reliability Diagram. If the reliability diagram is monotonic, it indicates that the average accuracy of low-confidence answers is always lower than that of high-confidence answers. 2) Confidence Evenness (CE) refers to the uniformity of the density of each bar in the reliability diagram. Our experimental results indicate that, even within the same dataset, there is a significant difference in the confidence of the answers provided by language models for different questions. This suggests that, after calibration, we could filter out some answers with low confidence to enhance the credibility of the model-generated answers. We summarize our main contributions as follows:
@@ 2.1 Logit-based Method
where $t$ is the sampling temperature of language models and $|\mathcal{A}|$ is the size of candidate answer set $\mathcal{A}$. Recent studies indicate that good calibration can be achieved by adjusting the temperature of RLHF-LMs Kadavath et al. (2022); He et al. (2023). However, temperature-scaling Guo et al. (2017) often requires higher temperatures, such as above 2.0 Kadavath et al. (2022), which might cause the outputs of the language models to become too random. When the probabilities for model-generated tokens are inaccessible, a straightforward solution is to deploy sampling and use the frequency of the sampled result to estimate the probability of generating this token. For instance, given a question $\mathcal{Q}$, we could sample $K$ times to acquire a set of answers $\mathcal{A}$ containing $N$ distinct answers, and each answer with an associated frequency $n_{i}$. The probability of the model generating answer $a_{i}$ can be estimated by $\frac{n_{i}}{K}$. Therefore, we could estimate the confidence of language models by $\mathcal{P}_{\mathrm{sampled}}(a_{i})$.
@@ 2.2 Verbalization-based Method
This method requires the model to have a strong ability to follow instructions and strong self-awareness (know whether it knows something or not Yin et al. (2023)). Accordingly, verbalized confidence can be a floating-point number between 0 and 1, i.e., ‘0.8’. And it can be linguistic expressions, i.e., ‘Almost Certain’, ‘About Even’, ‘Unlikely’. Although this method is quite easy to implement, we find various different LMs always tend to output some fixed high confidence expressions.
@@ 3 Methodology
Figure 3: Our Proposed Method.
@@ 3.4 Confidence Estimation
As demonstrated in Sections 3.2 and 3.3, we can estimate the model’s Uncertainty for a given question $\mathcal{Q}$ and the fidelity $\mathbf{F}(\cdot)$ among its different candidate answers. Therefore, the confidence of the model in its answer $a_{i}$ for question $\mathcal{Q}$ is defined as follows:
@@ 4 Experiments
To validate the effectiveness of our proposed method, we conducted experiments on different RLHF-LMs such as GPT-3.5-Turbo [^1] https://openai.com/chatgpt, LLaMA2-Chat Touvron et al. (2023) and Baichuan2-13B-Chat Yang et al. (2023a). To mitigate the influence of the sampling algorithm, unless specifically stated otherwise, we use hyper-parameters with a temperature of 1.0 and set top_p as 1.0.
@@ 4.1 Experimental Setting
Metrics.
We utilize multiple metrics to evaluate. We bin the predictions from the model by their confidence and report the $\mathrm{ECE}$ (expected calibration error).
$\mathrm{ECE}_{M}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})|,$ (8)
where $M$ is the bin number of the reliability diagram, $N$ is the overall sample number, and $|B_{m}|$ is the number of samples located in $B_{m}$. This paper also defines two novel metrics to evaluate the calibration. The first one is $\mathrm{IPR}$ (Inverse Pair Ratio), which is used to measure the monotonicity of the reliability diagram. If the reliability diagram is monotonic, it indicates that the average accuracy of answers with low confidence is lower than the average accuracy of answers with high confidence.
$\mathrm{IPR}_{M}=\frac{\mathrm{IP}}{C_{K}^{2}},$ (9)
where $\mathrm{IP}$ is the inverse pair number in the reliable diagram, and $K$ is the bin number with a density larger than 0. Suppose the confidence of language models is always equal to their average accuracy on a specific dataset. In that case, we cannot filter out some answers with low confidence to improve the truthfulness of language models. Therefore, we suggest using the $\mathrm{CE}$ (Confidence Evenness) to evaluate the uniformity of the density of each bar in the reliability diagram.
@@
We compared our approach with different baselines for eliciting the confidence of language model. The Verb method involves prompting the model to output a floating-point number between 0 and 1 to represent its confidence immediately after providing an answerTian et al. (2023); Lin et al. (2022a). The Ling method entails having the language model express its confidence level in natural language Tian et al. (2023). Since commercial models like ChatGPT do not provide per-token logits, we employed a sampling technique to estimate the probability of token generation, referred to as the Sampled method. Unless otherwise specified, the Sampled method involves sampling 10 times. For open-source models like LLaMA2-Chat, we directly use the probability of token generation as the measure of the language model’s confidence, which we refer to as the Token method. All the prompt templates we use are shown in Appendix B.
@@ 5.2 Parameter-Scaling
5.3 Truly Well-Calibrated Confidence
Previous work mainly evaluates the calibration of language models through expected calibration error (ECE). This section will discuss the research question: “What Kind of Confidence is Truly Well-Calibrated?”. Figure 6 demonstrates the calibration of various methods. From the calibration perspective, we hope that the confidence and accuracy relationship is close to the curve $y=x$. Thus, we need to reduce the ECE by calibrating confidence. Meanwhile, we hope that the reliability diagram should be as monotonic as possible to ensure that the accuracy of the results generated with low confidence is lower than that of the results with high confidence. Therefore, we propose the Inverse Pair Ratio (IPR) to evaluate monotonicity. From the perspective of building a more honest system, we hope the model’s confidence should be more evenly distributed across different confidence intervals. For example, if a language model has an overall accuracy of 75% on the TruthfulQA dataset and the confidence of each question from the language model is always 75%, its ECE and IPR would be 0. In this case, we think that the confidence may not necessarily be a truly well-calibrated confidence because we could not exclude some low-confidence results based on the confidence from the language model. Although the prior distribution of the model’s confidence is unknown, our confidence estimation method finds that language models have different confidence for different questions and different answers to the same question. Thus, we propose a metric called Confidence Evenness (CE) to measure the distance between the posterior confidence distribution and the uniform distribution. We believe ECE, IPR, and CE evaluate calibration from different perspectives and there is a trade-off between these three metrics. We suggest that truly well-calibrated confidence should achieve a balance among these three metrics, ECE, IPR, and CE, rather than over-optimizing any of them.
@@ 6 Conclusion
In this paper, we decompose the language model confidence into the Uncertainty about the question and the Fidelity to the answer generated by language models. Through the decomposition, we propose a plug-and-play method to calibrate the confidence of language models. Through experiments with 6 RLHF-LMs on 4 multiple-choice question answering benchmarks, our method exhibits good calibration. Besides, we propose two novel metrics, $\mathrm{IPR}$ and $\mathrm{CE}$, to evaluate the calibration of language models. Finally, we conduct a detailed discussion on Truly Well-Calibrated Confidence. We believe our method can serve as a strong baseline, and we hope that this work will provide some insights into the language model confidence calibration.
@@
Appendix A Algorithm
@@
Appendix B Prompt Templates
Linguistic Expression Confidence Score
‘Certain’ 1.0
‘Almost Certain’ 0.95
‘Highly Likely’ 0.9
‘Very Good Chance’ 0.8
‘We Believe’ 0.75
‘Probably’ 0.7
‘Probable’ 0.7
‘Likely’ 0.7
‘Better than Even’ 0.6
‘About Even’ 0.5
‘Probably Not’ 0.25
‘We Doubt’ 0.2
‘Unlikely’ 0.2
‘Little Chance’ 0.1
‘Chances are Slight’ 0.1
‘Improbable’ 0.1
‘Highly Unlikely’ 0.05
‘Almost No Chance’ 0.02
‘Impossible’ 0.0
Table 5: The EXPRESSION_LIST we used for the Ling Method.
Appendix C Why could CE be used as a metric?
As mentioned in section  4.2, we found that Language models tend to prefer outputting expressions of certain confidence, such as ’Highly Likely’, 0.8, and 0.9. In the table 8, we have counted the occurrence of different confidence levels for various models on different datasets to demonstrate the model’s preference for certain confidence levels when using the Verb and Ling method.
Appendix D The Computation Cost of Eliciting Fidelity
Model ARC-Challenge MMLU CommonSenseQA TruthfulQA Avg.
GPT-3.5-Turbo 2.774 2.984 3.052 3.275 3.021
GPT-4-Turbo 1.492 1.915 2.157 1.616 1.795
Baichuan2-13B-Chat 2.830 2.820 2.889 4.345 3.221
Llama2-7B-Chat 2.467 2.631 2.771 3.944 2.953
Llama2-13B-Chat 2.725 2.875 2.956 4.100 3.164
Llama2-70B-Chat 2.384 2.563 2.455 3.284 2.671
Appendix E Reliability Diagram
We provide the reliability diagrams of all the RLHF-LMs we evaluated in Figures 8-13. In a reliability diagram, the darker the color of the bar, the greater its density is, which indicates a preference for the confidence the language models express. Although the average accuracy of various RLHF-LMs is quite different, these model always prefer to express their confidence about 70-90% in verbalized methods.
Appendix F Additional Results
$\mathrm{BrierScore}_{M}=\frac{1}{N}\sum_{t=1}^{N}(f_{t}-o_{t})^{2},$ (11)
Method Prompt Template
[…]
Table 7: Prompt templates for each method evaluated.

### v2

@@ Abstract
Large language models optimized with techniques like RLHF have achieved good alignment in being helpful and harmless. However, post-alignment, these language models often exhibit overconfidence, where the expressed confidence does not accurately calibrate with their correctness rate. In this paper, we decompose the language model confidence into the Uncertainty about the question and the Fidelity to the answer generated by language models. Then, we propose a plug-and-play method, UF Calibration, to estimate the confidence of language models. Our method has shown good calibration performance by conducting experiments with 6 RLHF-LMs on four MCQA datasets. Moreover, we propose two novel metrics, IPR and CE, to evaluate the calibration of the model, and we have conducted a detailed discussion on Truly Well-Calibrated Confidence for large language models. Our method could serve as a strong baseline, and we hope that this work will provide some insights into the model confidence calibration.
@@
Mozhi Zhang1,2thanks: Work done during internship at Meituan., Mianqiu Huang1,2, Rundong Shi3, Linsen Guo 3, Chong Peng 3, Peng Yan 3, Yaqian Zhou1,2 Xipeng Qiu1,2thanks: Corresponding author. 1School of Computer Science, Fudan University 2Shanghai Key Laboratory of Intelligent Information Processing, Fudan University 3 Meituan {mzzhang22, mqhuang23}@m.fudan.edu.cn {shirundong}@meituan.com {zhouyaqian, xpqiu}@fudan.edu.cn
@@ 1 Introduction
Large language models (LLMs) acquire vast world knowledge and demonstrate powerful capabilities through pre-training Brown et al. (2020); OpenAI (2023); Bubeck et al. (2023); Sun et al. (2024). With technologies like RLHF Ouyang et al. (2022) and RLAIF Bai et al. (2022); Lee et al. (2023), large language models can become more helpful and harmless to align with human preferences Askell et al. (2021). However, how to build a more honest system has not yet been fully discussed. An honest model should have a certain understanding of the boundary of its knowledge, that is, knowing what it does not know Yin et al. (2023); Yang et al. (2023b); Zhou et al. (2024). A plausible method is utilizing the calibrated confidence to estimate the knowledge boundary of language models. For pre-trained language models, the per-token logit can already be considered a well-calibrated confidence score, which implies that pre-trained language models (mostly) know what they know Kadavath et al. (2022).
@@
Figure 1: In four different MCQA datasets, our method has demonstrated good calibration effects, meaning it is sufficiently close to the $y=x$ curve. The experimental data is derived from GPT-3.5-Turbo.
@@
As shown in Figure 2 and Appendix Tabel 6, by replacing the model’s answer with “All other options are wrong.”, we can assess whether the model had high fidelity to its previously given answer. Inspired by this phenomenon, we decompose the language model confidence into two dimensions: the Uncertainty about the question and the Fidelity to the answer generated by language models. First, if the answers provided by language model are consistent under multiple samplings, it indicates that language model has lower uncertainty regarding that question. Thus, we could utilize the information entropy of the frequency distribution of sampled answers to calculate the model’s uncertainty about a question. Second, we design a novel method to estimate the model’s fidelity to each of its sampled answers. Last, the uncertainty regarding question $\mathcal{Q}$ and the fidelity to the answer $a_{i}$ together determine the model’s confidence. As shown in Figure 1, our proposed UF Calibration achieved good calibration across different MCQA datasets. Meanwhile, UF Calibration does not require knowledge of the model’s per-token log-probability, making it broadly applicable to various Black-box RLHF-LMs, which do not provide the per-token log-probability.
To have a closer look at the calibration of model confidence, we propose two novel metrics for evaluating and observation: 1) Inverse Pair Ratio (IPR), which is the proportion of inverse pairs in the Reliability Diagram. This metric could reflect whether the model is well-calibrated from the perspective of the monotonicity of the Reliability Diagram. If the reliability diagram is monotonic, it indicates that the average accuracy of low-confidence answers is always lower than that of high-confidence answers. 2) As shown in Table 7, we find that as the number of model parameters increases, language models still tend to consistently express uncertainty within certain fixed ranges. Thus, we design the Confidence Evenness (CE) to observe to the uniformity of the density of each bar in the reliability diagram. Our experimental results indicate that, after calibration, even within the same dataset, there is a significant difference in the confidence of the answers provided by language models for different questions. We summarize our main contributions as follows:
@@ 2.1 Logit-based Method
where $t$ is the sampling temperature of language models and $|\mathcal{A}|$ is the size of candidate answer set $\mathcal{A}$. Recent studies indicate that good calibration can be achieved by adjusting the temperature of RLHF-LMs Kadavath et al. (2022); He et al. (2023). However, temperature-scaling Guo et al. (2017) often requires higher temperatures, such as above 2.0 Kadavath et al. (2022), which might cause the outputs of the language models to become too random. When the probabilities for model-generated tokens are inaccessible, a straightforward solution is to deploy sampling and use the frequency of the sampled result to estimate the probability of generating this token. For instance, given a question $\mathcal{Q}$, we could sample $K$ times to acquire a set of answers $\mathcal{A}$ containing $N$ distinct answers, and each answer with an associated frequency $n_{i}$. The probability of the model generating answer $a_{i}$ can be estimated by $\frac{n_{i}}{K}$. Therefore, we could estimate the confidence of language models by $\mathcal{P}_{\mathrm{sampled}}(a_{i})$. Recently,  Kumar et al. (2023) also propose to utilize the conformal prediction to calibrate the confidence of LLMs.
@@ 2.2 Verbalization-based Method
This method requires the model to have a strong ability to follow instructions and strong self-awareness (know whether it knows something or not Yin et al. (2023)). Accordingly, verbalized confidence can be a floating-point number between 0 and 1, i.e., ‘0.8’. And it can be linguistic expressions, i.e., ‘Almost Certain’, ‘About Even’, ‘Unlikely’. Although this method is quite easy to implement, we find various different LMs always tend to output some fixed high confidence expressions, as show in Table 7.
@@ 3 Methodology
Figure 3: Our proposed UF Calibration, which requires at most two phases to invoke the model. In the Sampling phase, for black-box models, similar to the Sampled method, we need to sample 10 times. For white-box models, a single invocation is sufficient. In the eliciting the fidelity phase, the model needs to be invoked approximately 2 to 3 times to generate a fidelity chain, as show in Table 5.
@@ 3.4 Confidence Estimation
Given the model’s Uncertainty for a given question $\mathcal{Q}$ and the fidelity $\mathbf{F}(\cdot)$ among different candidate answers, the confidence of the model in its answer $a_{i}$ for question $\mathcal{Q}$ is defined as follows:
@@ 4 Experiments
To validate the effectiveness of our proposed method, we conducted experiments on different RLHF-LMs such as GPT-3.5-Turbo [^1] https://openai.com/chatgpt, GPT-4-Turbo OpenAI (2023), LLaMA2-Chat Touvron et al. (2023) and Baichuan2-13B-Chat Yang et al. (2023a). To mitigate the influence of the sampling algorithm, unless specifically stated otherwise, we use hyper-parameters with a temperature of 1.0 and set top_p as 1.0.
@@ 4.1 Experimental Setting
Metrics.
We utilize multiple metrics to evaluate. We bin the predictions from the model by their confidence and report the $\mathrm{ECE}$ (expected calibration error). We also report the Brier Score of different methods in Table 11. In this paper, we also defines two novel metrics to evaluate the calibration. The first one is $\mathrm{IPR}$ (Inverse Pair Ratio), which is used to measure the monotonicity of the reliability diagram. If the reliability diagram is monotonic, it indicates that the average accuracy of answers with low confidence is lower than the average accuracy of answers with high confidence.
$\mathrm{IPR}_{M}=\frac{\mathrm{IP}}{C_{K}^{2}},$ (8)
where $\mathrm{IP}$ is the inverse pair number in the reliable diagram, and $K$ is the bin number with a density larger than 0. We found that as the number of model parameters increases, the accuracy of the model improves across various datasets. However, language models still tend to consistently express uncertainty within certain fixed ranges, and ECE cannot clearly reflect this phenomenon. Therefore, we suggest using the $\mathrm{CE}$ (Confidence Evenness) to evaluate the uniformity of the density of each bar in the reliability diagram.
@@
We compared our approach with different baselines for eliciting the confidence of language model. First, we reproduced the Verb and Ling method proposed by Tian et al. (2023). The Verb method involves prompting the model to output a floating-point number between 0 and 1 to represent its confidence immediately after providing an answer Tian et al. (2023); Lin et al. (2022a). The Ling method entails having the language model express its confidence level in natural language Tian et al. (2023). Since commercial models like ChatGPT do not provide per-token logits, we employed a sampling technique to estimate the probability of token generation, referred to as the Sampled method. Unless otherwise specified, the Sampled method involves sampling 10 times. For open-source models like LLaMA2-Chat, we directly use the probability of token generation as the measure of the language model’s confidence, which we refer to as the Token method. We also compare the Conformal Prediction Baseline proposed by Kumar et al. (2023) with our UF calibration in Appendix D.1. All the prompt templates we use are shown in Appendix E.
@@ 5 Analysis and Discussion
Truly Well-Calibrated Confidence
Previous work mainly evaluates the calibration of language models through ECE. This section will discuss the research question: “What Kind of Confidence is Truly Well-Calibrated?”. Figure 6 demonstrates the calibration of various methods. From the calibration perspective, we hope that the confidence and accuracy relationship is close to the curve $y=x$. Thus, we need to reduce the ECE by calibrating confidence. Meanwhile, we hope that the reliability diagram should be as monotonic as possible to ensure that the accuracy of the results generated with low confidence is lower than that of the results with high confidence. Therefore, we propose the Inverse Pair Ratio (IPR) to evaluate monotonicity. From the perspective of building a more honest system, we hope the model’s confidence should be distributed across different confidence intervals. For example, if a language model has an overall accuracy of 75% on the TruthfulQA dataset and the confidence of each question from the language model is always 75%, its ECE and IPR would be 0. And we find that different models tend to express confidence within a fixed interval. In this case, we think that the confidence may not necessarily be a truly well-calibrated confidence because we could not exclude some low-confidence results based on the confidence from the language model. Although the prior distribution of the model’s confidence is unknown, our confidence estimation method finds that language models have different confidence for different questions. Thus, we propose a metric called Confidence Evenness (CE) to measure whether the model confidence always is located in a fixed interval. We believe ECE, IPR, and CE evaluate calibration from different perspectives and there is a trade-off between these three metrics. We suggest that truly well-calibrated confidence should achieve a balance among ECE, IPR, and CE, rather than over-optimizing any of them.
@@ 6 Conclusion
In this paper, we decompose the language model confidence into the Uncertainty about the question and the Fidelity to the answer generated by language models. Through the decomposition, we propose a plug-and-play method, UF Calibration, to calibrate the confidence of language models. Through experiments with 6 RLHF-LMs on 4 multiple-choice question answering benchmarks, our method exhibits good calibration. Besides, we propose two novel metrics, $\mathrm{IPR}$ and $\mathrm{CE}$, to evaluate the calibration of language models. Finally, we conduct a detailed discussion on Truly Well-Calibrated Confidence. We believe our method can serve as a strong baseline, and we hope that this work could provide some insights into the language model confidence calibration.
@@
Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 62441602). The computations in this research were performed using the CFFF platform of Fudan University.
Appendix A The Computation Cost of Eliciting Fidelity
Model ARC-Challenge MMLU CommonSenseQA TruthfulQA Avg.
GPT-3.5-Turbo 2.774 2.984 3.052 3.275 3.021
GPT-4-Turbo 1.492 1.915 2.157 1.616 1.795
Baichuan2-13B-Chat 2.830 2.820 2.889 4.345 3.221
Llama2-7B-Chat 2.467 2.631 2.771 3.944 2.953
Llama2-13B-Chat 2.725 2.875 2.956 4.100 3.164
Llama2-70B-Chat 2.384 2.563 2.455 3.284 2.671
Appendix B Algorithm
@@
Model Is the answer chosen in the first round correct? Choose "All other options are wrong." after replacing Do not choose "All other options are wrong." after replacing
GPT-3.5-Turbo True 25.99% 33.27%
False 5.85% 34.88%
Acc. 81.61% 48.82%
GPT-4-Turbo True 70.75% 16.83%
False 3.00% 9.42%
Acc. 95.93% 64.10%
Baicuan2-13B-Chat True 5.14% 29.40%
False 4.22% 61.24%
Acc. 54.90% 32.43%
Llama2-7B-Chat True 3.92% 23.50%
False 4.83% 67.75%
Acc. 44.76% 25.75%
Llama2-13B-Chat True 3.55% 25.64%
False 2.82% 67.99%
Acc. 55.77% 27.39%
Llama2-70B-Chat True 13.59% 38.43%
False 3.98% 44.00%
Acc. 77.35% 46.62%
Table 6: We found that if the option chosen by the model in the first round is replaced with "All other options are wrong," the model then chooses "All other options are wrong" in the second round. In this case, the accuracy of the model’s first-round choice is significantly higher compared to when it chooses other options in the second round. The results are derived from TruthfulQA.